OpenStudy (anonymous):

differentiate f(x)=x/(x+(5/x))

OpenStudy (anonymous):

f(x)=x/(x+(5/x)) = (x^2/(x^2+5) ===> f'(x) = (2x^3+2x^2+10)/(x^2+5)^2